OpenStudy (anonymous):

8|5w-1|=0 and 12-+|3r+2| how about these one I really don't have an idea.

OpenStudy (anonymous):

\[8|5w-1|=0\iff 5w-1=0\] solve for \(w\)

OpenStudy (anonymous):

i don't understand the second one, my guess is it is a typo

OpenStudy (anonymous):

Can u answer the first one so i can do the second coz i really dnt have a idea

OpenStudy (raffle_snaffle):

so you are solving for w?

OpenStudy (anonymous):

\[5w-1=0\iff 5w=1\iff 5w=\frac{1}{5}\]

OpenStudy (anonymous):

Is that the answerr of that i dont get it wheres the 8

OpenStudy (nincompoop):

@captainmico88 coocoobird, make sure you copy the correct question and repost